# 2d Conformal Field Theories on Magic Triangle

### v1

@@ 3.2 Magic coset
We find that the coset relation (28) within the Cvitanovi’c–Deligne exceptional series is in fact a very special instance of a much broader class of coset constructions that exist among all theories in the extended magic triangle. Remarkably, for any two theories in the triangle one may form a coset that yields a third theory, which again belongs to the same triangle. More precisely, we uncover the following universal coset relation valid throughout the extended magic triangle:
@@
both of which were previously observed in Lee et al. (2024b). Equation (48) indicates that by taking cosets among exceptional series one can generate every theory in the extended magic triangle at level one.
@@
This form is particularly useful, as it implies that every theory in the extended magic triangle can be realized as a coset between two theories belonging to the Cvitanović–Deligne exceptional series. In this sense, the magic coset construction provides a CFT analogue of the centralizer construction of Deligne–Gross for the original magic triangle Deligne and Gross (2002).
Recall the degeneracy of non-vacuum primary of $\mathcal{T}^{(5,\nu)}$ theory is denoted as $\mathcal{N}_{(\nu)}$, which satisfies $\mathcal{N}_{(\nu)}=\mathcal{N}_{(1/\nu)}.$ The coset (52) implies that the degeneracies of $\mathcal{T}^{(\mu,\nu)}$ theory are
$\mathcal{M}^{(\mu,\nu)}_{1,2,3}=\left(\mathcal{N}_{(\mu)}\mathcal{N}_{(\nu)},\mathcal{N}_{(\nu)},\mathcal{N}_{(\mu)}\right).$ (53)
@@
The magic coset (52) further determines the modular $S$-transformation of $\mathcal{T}^{(\mu,\nu)}$. While one might naively expect the solutions of the fourth-order MLDE to form a degree-four vector-valued modular form, it is more natural in the present context to regard the four characters as forming a $2\times 2$ matrix-valued modular form. [^2] We thank Don Zagier for suggesting the interpretation in terms of matrix-valued modular forms. Let us define the $2\times 2$ matrix $X^{(\mu,\nu)}$ for the characters $\chi_{i}$ of $\mathcal{T}^{(\mu,\nu)}$ theory as
$X^{(\mu,\nu)}=\begin{pmatrix}\chi_{0}&\chi_{2}\\ \chi_{3}&\chi_{1}\end{pmatrix}$ (54)
@@
where $S_{(\nu)}$ denotes the $2\times 2$ $S$-matrix of the theory $\mathcal{T}^{(5,\nu)}$ in the Cvitanović–Deligne exceptional series. Taking the determinant of (55) and using $\det S_{(\nu)}=-1$, we see that $\det(X^{(\mu,\nu)})=\chi_{0}\chi_{1}-\chi_{2}\chi_{3}$ is modular invariant. Moreover, since $\alpha_{0}+\alpha_{1}=0$ and $\alpha_{2}+\alpha_{3}=1$, this determinant is holomorphic and must therefore be constant. A direct computation shows that the constant equals $R_{1}$, thereby proving the magic bilinear relation (46).
@@
Extracting the leading terms yields the following relations among representation dimensions:
@@
Here $R_{(\nu)}$ denotes the dimension of the fundamental representation of the Cvitanović–Deligne exceptional series, cf. (2.3). These relations are readily checked to be consistent with (47). For the row $\mathcal{T}^{(5,\nu)}$, the expression for $R_{3}^{(\mu,\nu)}$ becomes linear in $\nu$ and takes the form
$E_{7\frac{1}{2}}:\ 12\nu-3,\ E_{7}:12\nu-4,\ E_{6}:6\nu-3,\ F_{4}:6\nu-4$ (71)
$D_{4}:2\nu-2,\ G_{2}:2\nu-3,\ A_{2}:\nu-2,\ A_{1}:\nu-3,\ A_{\frac{1}{2}}:\nu-4.$ (72)
@@ 3.3 Five atomic theories
An interesting consequence of the magic coset relation (48) is that all thirty CFTs appearing in the extended magic triangle in Table 2 can be generated from only five atomic theories lying below the line $\mu\nu=1$. These five atoms are
$\mathrm{Vir}_{5,2}^{\rm eff},\ \mathrm{Vir}_{5,3}^{\rm eff},\ U(1)_{3},\ \mathrm{Vir}_{6,5}^{\rm e},\ D_{2\rm A}.$ (80)
This means that every theory $\mathcal{T}^{(\mu,\nu)}$ can be realized as an extension of a product of these atomic theories, or equivalently as a non-diagonal modular invariant of an tensor product CFT. More concretely, given a theory $\mathcal{T}^{(\mu,\nu)}$ in Table 2, we draw a right triangle bounded by the three lines $\mu=\mathrm{const}$, $\nu=\mathrm{const}$, and $\mu\nu=1$. The theory $\mathcal{T}^{(\mu,\nu)}$ is then constructed from the atomic theories that lie along the segment of the slope just below the line $\mu\nu=1$ within this triangle. Since each of the five atoms appears twice along this line, any theory in the extended magic triangle involves at most two identical atomic constituents. The basic data of these five atoms are summarized in Table 4.
@@ 3.4 The diagonal exceptional series
This relation provides a direct generalization of the $E_{8}$ dual pairs appearing in (28). For instance, when $\mu=3$, equation (81) implies that the RCFTs in the subexceptional series at level one naturally pair via coset constructions with respect to $(D_{6})_{1}$. Given the special interest on $\mathcal{T}^{(\mu,\mu)}$ theories which are
@@
we call them the diagonal exceptional series. Each of these theories admits a $\mathbb{Z}_{2}$ inner automorphism such that $h_{2}=h_{3}$. Consequently, the theory $\mathcal{T}^{(\mu,\mu)}$ possesses three independent characters with conformal weights $h_{i}=0,$ $(\mu-1)/(\mu+1)$, $\mu/(\mu+1)$. These characters satisfy a third-order holomorphic modular linear differential equation of the form
@@ 4 Magic triangle at level two
At level two, the magic triangle no longer behaves as perfect as at level one. For the 18 elements belonging to the original magic triangle in Table 1, the associated higher-level theories present no conceptual difficulties, and the corresponding RCFTs are well defined. By contrast, the remaining 12 theories in the extended triangle require more careful treatment at higher levels. On the one hand, the two elements with vanishing dimension $d=0$ and dual Coxeter number $h^{\vee}=-1$ become trivial at higher levels, since their central charges vanish identically. On the other hand, some intermediate Lie algebras do not admit consistent higher-level generalizations. A typical example is the $E_{7\frac{1}{2}}$ case, for which we have argued that only the levels $k\leq 5$ can be realized as RCFTs Lee et al. (2024a).
@@
$c_{2}=\frac{24(\mu\nu-1)(\mu\nu+\mu+\nu-4)}{(\mu+1)(\nu+1)(\mu\nu+\mu+\nu-3)}$ (88)
In principle, all 13 conformal weights can be written explicitly as functions of $\mu$ and $\nu$. In particular, there is always a vacuum character with exponent zero. We expect that the characters of all level-two theories satisfy a two-parameter family of holomorphic modular linear differential equations of order 13 with index $l=36$. Owing to their complexity, we do not present these equations explicitly. Nevertheless, substantial simplifications occur for certain exceptional series. For example, the characters of the CD exceptional series at level two satisfy a one-parameter family of sixth-order holomorphic MLDEs Lee et al. (2024a). Moreover, the intermediate exceptional series at level two are realized by minimal $\mathcal{W}$-algebras $\mathcal{W}_{-h^{\vee}/6+1}(\mathfrak{g}_{\rm CD},f_{\theta})$, which has been studied recently in Sun (2024); Arakawa et al. (2024); Kac and Wakimoto (2025). Finally, for the subexceptional series at level two, an emergent supersymmetry plays a crucial role. In this case, it is most natural to fermionize the theories and study their fermionic characters and fermionic MLDEs. A detailed analysis of this structure will be presented in the next section. Finally, we expect that the Severi exceptional series at level two should admit a $\mathbb{Z}_{3}$ parafermionic structure. It would be interesting to investigate this in the future.
@@ 4.1 Subexceptional series at level two
A remarkable phenomenon we observe is that all theories in subexceptional series $(\mu,\nu)=(3,\nu)$ at level two are supersymmetric. These theories $\mathcal{T}_{2}^{(3,\nu)}$ are
@@
A prerequisite of emergent supersymmetry is that the RCFT has a primary field with conformal weight $3/2$, which serves as the supersymmetry generator. It is easy to verify this is true for all theories in (88). In many cases, the supersymmetry of these theories has been studied, see e.g. Johnson-Freyd (2021); Bae and Lee (2021). In particular, the $(D_{6+1/2})_{2}$ was found to be supersymmetric recently in Lee et al. (2024b). Taking $\mu=3$ in (88), the central charge of $\mathcal{T}_{2}^{(3,\nu)}$ theory is
@@
We observe that the fermionzation of each theory in (89) denoted as $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ has at most four distinct Neveu–Schwarz characters and equal number of Ramond characters. In general, we find the four Neveu–Schwarz weights are
@@
while the four Ramond weights are
@@
Notice the first Ramond weight has exponent 0 which indicates the existence of the constant $\widetilde{\rm R}$ character. One can also easily recover 13 bosonic conformal weights and check that the 13 conformal weights always yield a 13th order MLDE with index $l=36$. For theories with less than four distinct NS characters, the fermionic weights take value in the above list.
We further find a one-parameter family of holomorphic fermionic MLDE of degree four for the Neveu–Schwarz characters of $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ theories. To present the fermionic MLDE, we first recall the Jacobi theta functions
@@
In the meantime, the Ramond characters should satisfy the following dual MLDE
@@
The eight fermionic weights (91) and (92) along with the no free fermion condition allow us to solve the unfixed coefficients $\lambda_{i}$ are rational functions of $\nu$. In the end, we find the one-parameter family as
@@
and $\lambda_{8}=2\lambda_{7}-2\lambda_{6}$. The last equation comes from the supersymmetry condition that one Ramond weight has zero exponent. We checked the Neveu–Schwarz and Ramond characters of $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ theories indeed satisfy the above fermionic MLDEs. This one-parameter family of fermionic MLDE suggests that all Neveu–Schwarz and Ramond characters of $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ theory can be solved as functions of $\nu$ in a uniform way.
@@ 4.2 Coset constructions
At level two, there no longer exists a universal coset construction covering the entire magic triangle, in contrast to the level-one case. Nevertheless, the magic coset relation (48) continues to provide strong guidance and motivates us to search for new uniform coset constructions involving the Cvitanović–Deligne, sub and Severi exceptional series at level two.
First we find the following universal coset construction between Cvitanović–Deligne and sub exceptional series at level two:
@@
Each of the four characters of $(G_{2})_{2}$ decomposes to the summation of 12 terms of $(A_{1})_{6}\times(A_{1})_{2}\times\mathrm{Vir}_{9,8}$. For instance, denote $\mathrm{Vir}_{9,8}$ as $M$, then for the vacuum we have
@@
We further find the following universal coset construction between the sub and Severi exceptional series at level 2:
@@
Here again $\nu=\frac{1}{2},\frac{2}{3},\frac{3}{2},2,3,5$. The $\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1}$ is the bosonization $\mathcal{B}$ of the unitary $N=1$ supersymmetric minimal model $\mathrm{Vir}_{n+2,n}^{N=1}$. For example, taking $\nu=3$ gives
@@
It is easy to find $\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ has in total $18$ bosonic primary fields with conformal weights
@@
Denote $\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ as $\mathcal{B}$, then for the vacuum, we find the character character relation
@@
We check the degeneracy matches and it is straightforward to upgrade the character relation to module decomposition.
Furthermore, there exists the following universal coset Cheng et al. (2020) for Cvitanović–Deligne exceptional series resembling the Maverick coset
@@
Here $\nu$ takes value in $1/3,1/2,2/3,3/2,2,3,5$. For $\nu=1/4,4$, it still holds if we scale $n$ and $n+1$ such that they are coprime integers and take the proper non-diagonal modular invariant. For $\nu=4$, the non-diagonal modular invariant was identified in Lee et al. (2024a). For $\nu=1/4$, the coset was recently proved in Feng and Jiang (2025). We newly find a similar coset construction for the subexceptional series:
$\frac{(\mathfrak{g}_{\rm sub})_{1}\times(\mathfrak{g}_{\rm sub})_{1}}{(\mathfrak{g}_{\rm sub})_{2}}=\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1},\qquad n=\left\{\begin{array}[]{ll}2+\frac{4}{\nu-1},&\nu>1,\\ \frac{4}{1-\nu}-4,&\nu<1.\end{array}\right$. (112)
We have a final remark for the intermediate exceptional series at level 2. For $\nu=2,3,5$ and $n=3+\frac{4}{\nu-1}$, the following cosets were found in (Sun, 2024, Equation (1.9),(1.10)):
$\frac{\mathcal{T}^{(5,\nu)}_{2}}{\mathcal{T}^{(4,\nu)}_{2}}$ $=\frac{(\mathfrak{g}_{\rm CD})_{2}}{(\mathfrak{g}_{I})_{2}}=(\mathrm{Vir}_{3n+1,n})_{\rm eff},$ (113)
$\frac{\mathcal{T}^{(4,\nu)}_{2}}{\mathcal{T}^{(3,\nu)}_{2}}$ $=\frac{(\mathfrak{g}_{I})_{2}}{(\mathfrak{g}_{\rm sub})_{2}}=(\mathcal{B}\mathrm{Vir}^{N=1}_{3n+1,n+1})_{\rm eff}.$ (114)

### v2

@@ 3.2 Magic coset
We find that the coset relation (28) within the Cvitanović–Deligne exceptional series is in fact a very special instance of a much broader class of coset constructions that exist among all theories in the extended magic triangle. Remarkably, for any two theories in the magic triangle one may form a coset that yields a third theory, which again belongs to the magic triangle. More precisely, we uncover the following universal coset relation valid throughout the extended magic triangle:
@@
both of which were previously observed in Lee et al. (2024b). Equation (48) also indicates that by taking cosets among exceptional series one can generate every theory in the extended magic triangle at level one.
@@
This form is particularly useful, as it implies that every theory in the extended magic triangle in Table 2 can be realized as a coset between two theories belonging to the Cvitanović–Deligne exceptional series. In this sense, the magic coset construction provides a CFT analogue of the centralizer construction of Deligne–Gross for the original magic triangle Deligne and Gross (2002).
Recall the degeneracy of the non-vacuum primary field of $\mathcal{T}^{(5,\nu)}$ theory is denoted as $\mathcal{N}_{(\nu)}$, which satisfies $\mathcal{N}_{(\nu)}=\mathcal{N}_{(1/\nu)}.$ The numbers are collected in Table 3. The coset (52) implies that the degeneracies of $\mathcal{T}^{(\mu,\nu)}$ theory are
$(\mathcal{M}^{(\mu,\nu)}_{1},\mathcal{M}^{(\mu,\nu)}_{2},\mathcal{M}^{(\mu,\nu)}_{3})=\left(\mathcal{N}_{(\mu)}\mathcal{N}_{(\nu)},\mathcal{N}_{(\nu)},\mathcal{N}_{(\mu)}\right).$ (53)
@@
The magic coset (52) further determines the modular $S$-transformation of $\mathcal{T}^{(\mu,\nu)}$. While one might expect the solutions of the fourth-order MLDE to form a degree-four vector-valued modular form, it is more natural in the present context to regard the four characters as forming a $2\times 2$ matrix-valued modular form. [^5] We thank Don Zagier for suggesting the interpretation in terms of matrix-valued modular forms. Let us define the $2\times 2$ matrix $X^{(\mu,\nu)}$ for the characters of $\mathcal{T}^{(\mu,\nu)}$ theory as
$X^{(\mu,\nu)}=\begin{pmatrix}\chi_{0}&\chi_{2}\\ \chi_{3}&\chi_{1}\end{pmatrix}$. (54)
@@
where $S_{(\nu)}$ denotes the $2\times 2$ $S$-matrix of the theory $\mathcal{T}^{(5,\nu)}$ in the Cvitanović–Deligne exceptional series, given explicitly in B. Taking the determinant of (55) and using the fact $\det S_{(\nu)}=-1$, we see that $\det(X^{(\mu,\nu)})=\chi_{0}\chi_{1}-\chi_{2}\chi_{3}$ is modular invariant. Moreover, since $\alpha_{0}+\alpha_{1}=0$ and $\alpha_{2}+\alpha_{3}=1$, this determinant is holomorphic and must therefore be constant. A direct computation shows that the constant equals $R_{1}$, thereby proving the magic bilinear relation (46).
@@
Extracting the leading terms in $q$-expansion yields the following relations among representation dimensions:
@@
Here $R_{(\nu)}$ denotes the dimension of the fundamental representation of the Cvitanović–Deligne exceptional series, cf. (2.3) and Table 3. These relations are readily checked to be consistent with (47). For the row $\mathcal{T}^{(5,\nu)}$, the expression for $R_{3}^{(\mu,\nu)}$ becomes linear in $\nu$ and takes the form
$E_{7+1/2}:12\nu-3,\quad E_{7}:12\nu-4,\quad E_{6}:6\nu-3,\quad F_{4}:6\nu-4,$ (71)
$D_{4}:2\nu-2,\quad G_{2}:2\nu-3,\quad A_{2}:\nu-2,\quad A_{1}:\nu-3,\quad A_{1/2}:\nu-4.$ (72)
@@ 3.3 Five atomic theories
An interesting consequence of the magic coset relation (48) is that all thirty CFTs appearing in the extended magic triangle in Table 2 can be generated from only five atomic theories lying just below the line $\mu\nu=1$. These five atoms are
$\mathrm{Vir}_{5,2}^{\rm eff},\ \mathrm{Vir}_{5,3}^{\rm eff},\ (U_{1})_{3},\ \mathrm{Vir}_{6,5}^{\rm e},\ D_{2\rm A}.$ (80)
This means that every theory $\mathcal{T}^{(\mu,\nu)}$ can be realized as an extension of a product of these atomic theories, or equivalently as a non-diagonal modular invariant of an tensor product CFT of the atoms. More concretely, given a theory $\mathcal{T}^{(\mu,\nu)}$ in Table 2, we draw a right triangle bounded by the three lines $\mu=\mathrm{const}$, $\nu=\mathrm{const}$, and $\mu\nu=1$. The theory $\mathcal{T}^{(\mu,\nu)}$ is then constructed from the atomic theories that lie along the segment of the slope just below the line $\mu\nu=1$ within this triangle. Since each of the five atoms appears twice just below the $\mu\nu=1$ line, any theory in the extended magic triangle involves at most two identical atomic constituents. The basic data of these five atoms are summarized in Table 4.
@@ 3.4 The diagonal exceptional series
This relation provides a direct generalization of the $(E_{8})_{1}$ dual pairs appearing in (28). For instance, when $\mu=3$, equation (81) implies that the RCFTs in the subexceptional series at level one naturally pair via coset constructions with respect to $(D_{6})_{1}$. Given the special interest on $\mathcal{T}^{(\mu,\mu)}$ theories, namely,
@@
we refer to them as the diagonal exceptional series. Each of these theories admits a $\mathbb{Z}_{2}$ inner automorphism such that $h_{2}=h_{3}$. Consequently, the theory $\mathcal{T}^{(\mu,\mu)}$ possesses three independent characters with conformal weights $h_{i}=0,$ $(\mu-1)/(\mu+1)$, $\mu/(\mu+1)$. These characters satisfy a third-order holomorphic modular linear differential equation of the form
@@ 4 Magic triangle at level two
At level two, the magic triangle no longer behaves as perfectly as at level one. For the 18 elements that belong to the original magic triangle in Table 1, the associated higher-level theories do not present conceptual difficulties, and the corresponding RCFTs are well defined. By contrast, the remaining 12 theories in the extended triangle require more careful treatment at higher levels. On the one hand, the two elements with vanishing dimension $d=0$ and dual Coxeter number $h^{\vee}=-1$ become trivial at higher levels, since their central charges vanish identically. On the other hand, some intermediate Lie algebras do not admit consistent higher-level generalizations. A typical example is the case $E_{7\frac{1}{2}}$, for which we have argued that only levels $k\leq 5$ can be realized as RCFTs Lee et al. (2024a).
@@
$c_{2}=\frac{24(\mu\nu-1)(\mu\nu+\mu+\nu-4)}{(\mu+1)(\nu+1)(\mu\nu+\mu+\nu-3)}.$ (88)
In principle, all 13 conformal weights can be written explicitly as functions of $\mu$ and $\nu$. In particular, there is always a character with exponent zero. We expect that the characters of all level-two theories satisfy a two-parameter family of holomorphic modular linear differential equations of order 13 with index $l=36$. Owing to their complexity, we do not present these equations explicitly. Nevertheless, substantial simplifications occur for certain exceptional series. For example, the characters of the CD exceptional series at level two satisfy a one-parameter family of sixth-order holomorphic MLDEs Lee et al. (2024a). Moreover, the intermediate exceptional series at level two are realized by minimal $\mathcal{W}$-algebras $\mathcal{W}_{-h^{\vee}/6+1}(\mathfrak{g}_{\rm CD},f_{\theta})$, which has been studied in Arakawa and Moreau (2018); Sun (2024); Arakawa et al. (2024); Kac and Wakimoto (2025). Finally, for the subexceptional series at level two, an emergent supersymmetry plays a crucial role. In this case, it is most natural to fermionize the theories and study their fermionic characters and fermionic MLDEs. A detailed analysis of this structure will be presented in the next subsection. Finally, we expect that the Severi exceptional series at level two should admit a $\mathbb{Z}_{3}$ parafermionic structure. It would be interesting to investigate this in the future.
@@ 4.1 Subexceptional series at level two
A remarkable phenomenon we observe is that all theories in subexceptional series at level two are supersymmetric. These theories $\mathcal{T}_{2}^{(3,\nu)}$ are
@@
A prerequisite of emergent supersymmetry is that the RCFT has a primary field with conformal weight $3/2$, which serves as the supersymmetry generator. It is easy to verify this is true for all theories in (88). In many cases, the supersymmetry of these theories has been studied extensively, see e.g. Johnson-Freyd (2021); Bae and Lee (2021). In particular, the $(D_{6+1/2})_{2}$ was found to be supersymmetric only recently in Lee et al. (2024b). setting $\mu=3$ in (88), the central charge of $\mathcal{T}_{2}^{(3,\nu)}$ theory becomes
@@
We observe that the fermionization of each theory in (89), denoted as $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$, has at most four distinct Neveu–Schwarz characters and an equal number of Ramond characters. In general, we find that the four Neveu–Schwarz conformal weights are
@@
while the four Ramond conformal weights are
@@
Notice the first Ramond weight has exponent 0 which indicates the existence of the constant $\widetilde{\rm R}$ character. One can also easily recover 13 bosonic conformal weights and verify that the 13 conformal weights always yield a 13th order MLDE with index $l=36$. For theories with less than four distinct NS characters, the fermionic weights take values in the list above.
We further find a one-parameter family of holomorphic fermionic MLDE of degree four for the Neveu–Schwarz characters of the theories $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$. To present the fermionic MLDE, we first recall the Jacobi theta functions
@@
In the meantime, the Ramond characters satisfy the following dual MLDE obtained from modular transformation
@@
The eight fermionic weights (91) and (92) along with the no free fermion condition allow us to solve the unfixed coefficients $\lambda_{i}$ as rational functions of $\nu$. In the end, we find the one-parameter family given by
@@
$\lambda_{8}=2\lambda_{7}-2\lambda_{6}.$
The last equation comes from the supersymmetry condition that one Ramond conformal weight has zero exponent. We have explicitly checked that the Neveu–Schwarz and Ramond characters of the theories $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ indeed satisfy the above fermionic MLDEs. This one-parameter family of fermionic MLDE suggests that all Neveu–Schwarz and Ramond characters of the theories $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ can be solved uniformly as functions of $\nu$.
@@ 4.2 Coset constructions
In contrast to the level-one case, there no longer exists a universal coset construction covering the entire magic triangle at level two. Nevertheless, the magic coset relation (48) continues to provide strong guidance and motivates us to search for new uniform coset constructions involving the Cvitanović–Deligne, subexceptional, and Severi exceptional series at level two.
We first find the following universal coset construction between the Cvitanović–Deligne and subexceptional series at level two:
@@
Each of the four characters of $(G_{2})_{2}$ decomposes into the sum of 12 terms of $(A_{1})_{6}\times(A_{1})_{2}\times\mathrm{Vir}_{9,8}$. For instance, denote $\mathrm{Vir}_{9,8}$ as $M$. For the vacuum character, we have
@@
We further find the following universal coset construction between the subexceptional and Severi exceptional series at level 2:
@@
Here again $\nu=\frac{1}{2},\frac{2}{3},\frac{3}{2},2,3,5$. The theory $\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1}$ denotes the bosonization $\mathcal{B}$ of the unitary $N=1$ supersymmetric minimal model $\mathrm{Vir}_{n+2,n}^{N=1}$. As an example for $\nu=3$ the coset (104) gives
@@
It is straightforward to verify that find $\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ has in total $18$ bosonic primary fields with conformal weights
@@
Denote $\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ as $\mathcal{B}$, the vacuum character satisfies the following relation:
@@
We have checked that the degeneracies match, and it is straightforward to upgrade the character relation to module decomposition.
Furthermore, there exists the following universal coset Cheng et al. (2020) for Cvitanović–Deligne exceptional series resembling the Maverick coset Dunbar and Joshi (1993)
@@
Here $\nu=\frac{1}{2},\frac{2}{3},\frac{3}{2},2,3,5$. For $\nu=\frac{1}{4},4$, the relation continues to hold provided one scales $n$ and $n+1$ so that they are coprime integers and choose the appropriate non-diagonal modular invariant. For $\nu=4$, such a non-diagonal modular invariant was identified in Lee et al. (2024a). For $\nu=\frac{1}{4}$, the corresponding coset was recently proved in Feng and Jiang (2025), with $\mathfrak{g}^{(5,1/2)}$ as $osp(1|2)$. We further find a similar coset construction for the subexceptional series:
$\frac{\mathcal{T}^{(3,\nu)}_{1}\times\mathcal{T}^{(3,\nu)}_{1}}{\mathcal{T}^{(3,\nu)}_{2}}=\frac{(\mathfrak{g}_{\rm sub})_{1}\times(\mathfrak{g}_{\rm sub})_{1}}{(\mathfrak{g}_{\rm sub})_{2}}=\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1},\qquad n=\left\{\begin{array}[]{ll}2+\frac{4}{\nu-1},&\nu>1,\\ \frac{4}{1-\nu}-4,&\nu<1.\end{array}\right$. (112)
Finally, we make a remark concerning the intermediate exceptional series at level 2. For $\nu=2,3,5$ and $n=3+\frac{4}{\nu-1}$, the following cosets were found in (Sun, 2024, Equation (1.9),(1.10)):
$\frac{\mathcal{T}^{(5,\nu)}_{2}}{\mathcal{T}^{(4,\nu)}_{2}}$ $=\frac{(\mathfrak{g}_{\rm CD})_{2}}{(\mathfrak{g}_{I})_{2}}=(\mathrm{Vir}_{3n+1,n})^{\rm eff},$ (113)
$\frac{\mathcal{T}^{(4,\nu)}_{2}}{\mathcal{T}^{(3,\nu)}_{2}}$ $=\frac{(\mathfrak{g}_{I})_{2}}{(\mathfrak{g}_{\rm sub})_{2}}=(\mathcal{B}\mathrm{Vir}^{N=1}_{3n+1,n+1})^{\rm eff}.$ (114)